OpenStudy (anonymous):

cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5

OpenStudy (anonymous):

Help Please!

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

@ilikemath50

OpenStudy (anonymous):

Did not learn cube roots yet, Can't help you there, sorry

OpenStudy (anonymous):

Dang

OpenStudy (anonymous):

do you know anyone

OpenStudy (anonymous):

I'll try to send someone else in, @precal @thomaster @iambatman

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem! :)

OpenStudy (anonymous):

Im new to this like I just started

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

Yeah, when I just started all I did was ask questions, but im pretty good at math so I was like hey! why not help people with their questions? I have gotten quite a few medals and fans now.. I could do better though, lol

OpenStudy (precal):

give me a moment to try to post the equation

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (precal):

\[\frac{ \sqrt[3]{5}\sqrt{5} }{ \sqrt[3]{5^5} }\] is this your problem?

OpenStudy (anonymous):

Yes

OpenStudy (precal):

ok lets rewrite it using fractions

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

here are my answer choices

OpenStudy (precal):

\[\frac{ 5^{1/3}5^{1/2} }{ 5^{5/3} }\]

OpenStudy (precal):

don't need your choices

OpenStudy (anonymous):

ok

OpenStudy (precal):

ok what is x times x?

OpenStudy (anonymous):

1

OpenStudy (precal):

no x times x is x^2 (x to the second power)

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

I understand that

OpenStudy (precal):

why is this true? because we added the powers so \[x^1(x^2)=x^3\]

OpenStudy (precal):

ok we do need the same base for this to be true, btw this is one of your laws of exponents

OpenStudy (precal):

so do the top of the fraction for me

OpenStudy (anonymous):

ok you have to add the exponents

OpenStudy (anonymous):

5 5/6

OpenStudy (anonymous):

because i did 1/3 = 2/6 then 1/2 =3/6 right? @precal

OpenStudy (precal):

yes it is 5 and 5/6

OpenStudy (anonymous):

Yay ok now the next part

OpenStudy (precal):

now this leads us to another law of exponent \[\frac{ x^3 }{ x }\] is what?

OpenStudy (precal):

\[\frac{ x^3 }{ x }=\frac{ xxx }{ x }=xx=x^2\]

OpenStudy (precal):

do you see the pattern for this law of exponent?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

so its 5 5/6 over 5 5/3

OpenStudy (precal):

\[\frac{ a^b }{ a^c }=a ^{b-c}\]

OpenStudy (anonymous):

Hold on i gtg fix my blinds be back

OpenStudy (precal):

ok

OpenStudy (precal):

let me know when to continue

OpenStudy (anonymous):

k im back @precal I had new windows put in and They took my blinds down

OpenStudy (precal):

no problem, I fixed mine the other days.....I would rather solve math problems than real world house problems.....

OpenStudy (precal):

ok so the bases are the same and we are going to subtract the exponents. Always do top exponent minus bottom exponent

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so its 5 5/6 minus 5 5/3

OpenStudy (precal):

\[\frac{ 5^{5/6} }{ 5^{5/3} }\]

OpenStudy (anonymous):

would I make it 10/6

OpenStudy (precal):

\[5^{\frac{ 5 }{ 6}-\frac{ 5 }{ 3 }}\]

OpenStudy (precal):

should be -5/6

OpenStudy (anonymous):

THANK YOU

OpenStudy (precal):

\[5^{-5/6}\]

OpenStudy (anonymous):

How can I like favorite you

OpenStudy (anonymous):

Im new

OpenStudy (precal):

well we are not done, leads us to another law of exponent,

OpenStudy (anonymous):

O ok

OpenStudy (precal):

mark response look for the blue rectangle that states "best response" this gives us medals and you can "fan me" if you want, I think you have to pick something on my name.....

OpenStudy (precal):

last step

OpenStudy (precal):

\[x ^{-2}=\frac{ 1 }{ x^2 }\]

OpenStudy (anonymous):

so we have 5 -5/6

OpenStudy (precal):

negative powers move things from numerator to denominator so x^-2 is a numerator and to get rid of the negative sign, we make it a denominator

OpenStudy (precal):

\[\frac{ x^3y ^{-2} }{ z ^{-4} }\]

OpenStudy (anonymous):

were did you get that?

OpenStudy (precal):

ok see how y has a negative power and z has a negative power I am making up an example to help you understand the last step

OpenStudy (anonymous):

Yes o ok

OpenStudy (precal):

\[\frac{ x^3y ^{-2} }{ z ^{-4} }=\frac{ x^3z^4 }{ y^2 }\]

OpenStudy (precal):

see how y went to the bottom and z went to the top

OpenStudy (anonymous):

ok yes

OpenStudy (precal):

computer is acting weird, I need to get a new mouse

OpenStudy (anonymous):

Ok lol

OpenStudy (anonymous):

So is 5 -5/6 the answer?

OpenStudy (anonymous):

Or is there another part

OpenStudy (precal):

sorry, my computer is acting weird. I had to switch out three mouses

OpenStudy (precal):

\[\frac{ 1 }{ 5^{5/6} }\]

OpenStudy (precal):

this is your final answer

OpenStudy (anonymous):

Its doesnt show

OpenStudy (precal):

sometimes the system gets rid when there are a lot of people on the site

OpenStudy (precal):

get weird I meant

OpenStudy (precal):

can you see it now

OpenStudy (anonymous):

O wait I just rembered that the properties of exponents shows that if you have a 1 over the answer that it can be made into a negative answer!!!! so 5 -5/6 is right!!

OpenStudy (jdoe0001):

\(\Large \bf { a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\\quad \\ a^{-\frac{{\color{blue} n}}{{\color{red} m}}} = \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}} \implies \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}} \\ \quad \\ \quad \\ \frac{ \sqrt[3]{5}\sqrt[2]{5} }{ \sqrt[3]{5^5} }\implies \cfrac{5^{\frac{1}{3}}\cdot 5^{\frac{1}{2}}}{5^{\frac{5}{3}}}\implies \cfrac{5^{\frac{1}{3}}\cdot 5^{\frac{1}{2}}}{1}\cdot \cfrac{1}{5^{\frac{5}{3}}} \\ \quad \\ 5^{\frac{1}{3}}\cdot 5^{\frac{1}{2}}\cdot 5^{-\frac{5}{3}}\implies 5^{\frac{1}{3}+\frac{1}{2}-\frac{5}{3}=\frac{4+6-20}{12}=\frac{-\cancel{ 10 }}{\cancel{ 12 }}=\frac{-5}{6}} \\ \quad \\ 5^{\frac{-5}{6}}\implies \cfrac{1}{5^{\frac{5}{6}}}\implies \cfrac{1}{\sqrt[6]{5^5}} }\)

OpenStudy (anonymous):

THANK YOU GUYS SO MUCH ILL TELL YOU IF I GET IT RIGHT!!!

OpenStudy (precal):

yes but it is best to get rid of negative exponents

OpenStudy (precal):

anytime

OpenStudy (anonymous):

I actually have a couple more ill post them

OpenStudy (precal):

sure, close this problem out and open a new thread please